# An Embedding Theorem for tractor bundles, and an application in conformal pseudo-Riemannian geometry

### v1

@@ 3. Embedding Theorem
$\iota$ ::: $\hat{M}\rightarrow\textrm{Mon}({\mathcal{S}},{\bf V})$
@@
Let $\mathbb{U}=\mbox{Mon}({\mathcal{S}},{\bf V})/P$ and $\bar{\iota}:M\rightarrow\mathbb{U}$ the induced map, which is also ${\textrm{Aut}}(M,\omega)$-equivariant. Note that if $g\in{\textrm{Aut}}(M,\omega)$ stabilizes $x$, then $\psi(g)$ stablizes $\bar{\iota}(x)$. Although $\mathbb{U}$ is not generally an algebraic variety, the Rosenlicht stratification [30] (see also [14, 2.2]) yields a decomposition
@@
with ${\textrm{Aut}}(M,\omega)$ acting by precomposition via $\psi$ on the first factor and trivially on the second factor, commuting with the $P$-action on the product. Considering the map $\iota\times K:\hat{M}\rightarrow\mathbb{U}$ given by $\hat{x}\mapsto(\iota_{\hat{x}},K(\hat{x}))$, the same proof with the Borel Density Theorem as above yields, for almost-every $x\in M$, not only the embedding of $(\mathcal{S},\psi(B))$ into $({\bf V},\rho(P))$ but also that $K(\hat{x})$ is fixed by $\check{B}_{\hat{x}}$; the conclusion that $\check{B}_{\hat{x}}$ contains the isotropy of $B$ at $x$ with respect to $\hat{x}$ moreover remains true.
@@
$D^{m}\kappa:\hat{x}\mapsto\left(\kappa(\hat{x}),D\kappa(\hat{x}),\ldots,D^{(m)}\kappa(\hat{x})\right)$
takes its values (see Sections 2.1.2, 2.1.4). Fixed points of $P$ in this module can give rise to local automorphisms by the Frobenius Theorem for Cartan geometries. More precisely, for $(M,\hat{M},\omega)$ real-analytic and $M$ closed, there is $m\geq 0$ such that for any $\hat{x}\in\hat{M}$, any $p\in P$ fixing $D^{m}\kappa(\hat{x})$ is the isotropy at $x=\pi(\hat{x})$ with respect to $\hat{x}$ of a unique $g\in{\textrm{Aut}}^{loc}(M,\omega)$ fixing $x$ (where uniqueness applies to any two such local automorphisms defined in the same neighborhood of $x$) [21], see also [27] for a smooth version. The following corollary of the Embedding Theorem is well known in the setting of infinitesimal automorphisms (see [21, Thm 5.3]).
@@ 4. Application in conformal geometry
The remainder of the paper is devoted to the proof of Theorem 1.2. The first and longer part of the proof is algebraic and culminates in Proposition 4.12. We set $G=\mbox{O}(p+1,q+1)$ and $H$ a group locally isomorphic to $\mbox{SU}(p^{\prime},q^{\prime})$, with $p<q$ and $p^{\prime}<q^{\prime}$ as in the theorem.
@@
If $p<q$ then the root system of ${\mathfrak{o}}(p+1,q+1)$ is of type $B_{p+1}$, while if $p=q$, then it is type $D_{p+1}$. If $p^{\prime}<q^{\prime}$, then the root system of $\mathfrak{su}(p^{\prime},q^{\prime})$ is of type $BC_{p^{\prime}}$, while if $p^{\prime}=q^{\prime}$, then it is of type $C_{p^{\prime}}$.
Let $\beta_{1},\ldots,\beta_{p+1}\in{\mathfrak{a}}^{*}$ be such that the positive roots of $\mathfrak{g}$ are
@@
The root system is $B_{p+1}$. The set $\{\beta_{k}\}$ are called the positive short roots, and their root spaces are each of dimension $q-p$. The remaining two subsets are the long roots, such that the $\{\beta_{i}-\beta_{j}\}$ are the initial roots of a string of positive roots, while the $\{\beta_{i}+\beta_{j}\}$ are the terminal roots of a string of positive roots. Either of these types of roots have $1$-dimensional root spaces. These strings are such that for any $X\in{\mathfrak{g}}_{\beta_{i}-\beta_{j}}$ and $Y\in{\mathfrak{g}}_{\beta_{j}}$ both nonzero, the brackets $[X,Y]\in{\mathfrak{g}}_{\beta_{i}}$ and $[[X,Y],Y]\in{\mathfrak{g}}_{\beta_{i}+\beta_{j}}$ are nonzero. The weights of the $P$-representation on $\mathfrak{g}/{\mathfrak{p}}$ are
@@
in particular, the conformal factor is given by $-2\beta_{1}$.
@@
For the first application of the Embedding Theorem, we consider a Borel subgroup $B_{H}<H$. Note that $B_{H}$ is solvable, algebraic, and discompact. As $B_{H}$ is amenable and $M$ is compact, there is a $B_{H}$-invariant probability measure $\mu$ on $M$. Let $\check{B}=\check{B}_{\hat{x}}<\mbox{Ad}_{\mathfrak{g}}P$ be the algebraic subgroup given by Theorem 3.1 applied to $B_{H}$ acting on the adjoint tractor bundle endowed with Čap’s connection as in Section 1.1.3. The parallel sections correspond to infinitesimal automorphisms, and we restrict to $\mathcal{S}=\mathfrak{h}$. Note that this version of the embedding theorem is exactly the one of [3].
@@ 4.1. Alignment of the intertwining epimorphism
The Lie algebra $\check{{\mathfrak{b}}}$ is a subalgebra of ${\mathfrak{p}}$, with adjoint action preserving the image $\iota(\mathfrak{h})$, such that the restriction to this image is a surjection onto $\iota\circ\mbox{ad}_{\mathfrak{h}}{\mathfrak{b}}_{H}\circ\iota^{-1}\cong{\mathfrak{b}}_{H}$. We will denote by $r$ the resulting Lie algebra epimorphism $\check{{\mathfrak{b}}}\rightarrow{\mathfrak{b}}_{H}$. We can focus on the solvable radical $\check{{\mathfrak{b}}}^{solv}$, as any semisimple Levi factor is in the kernel of $r$. Because it is algebraic, $\check{{\mathfrak{b}}}^{solv}$ is invariant under Jordan decomposition (see [25, Thm 4.3.3]). Since Ad $\mbox{Ad }B_{H}$ has no elliptic elements, any elliptic elements of $\check{{\mathfrak{b}}}^{solv}$ are in the kernel of $r$; more precisely, they belong to a compact torus of $\ker r$, which is a direct factor of $\check{{\mathfrak{b}}}^{solv}$ (see [25, Thm 4.4.7, Cor 4.4.12]), so we will simply exclude this factor and assume that all elements of $\check{{\mathfrak{b}}}^{solv}$ have trivial elliptic components in their Jordan decomposition.
@@
Lie’s Theorem ensures that in $V^{p,q}$ there is a simultaneous eigenvector $v_{1}$ or a $2$-dimensional irreducible subspace for $\check{B}^{solv}$. In the latter case, the restriction of $\check{B}^{solv}$ splits into a nontrivial elliptic and a scalar action. By the assumption that all elements of $\check{B}^{solv}$ have trivial elliptic Jordan components, this possibility is excluded. If the eigenvector $v_{1}$ is timelike or spacelike, then the eigenvalues on it are ±$\pm 1$. In this case, apply Lie’s Theorem again to $v_{1}^{\perp}$, to conclude that it contains an isotropic simultaneous eigenvector, or $\check{{\mathfrak{b}}}^{solv}$ is trivial on $V^{p,q}$. Proceeding inductively yields a null-orthonormal basis in which $\check{B}^{solv}$ is upper-triangular. It follows that on $V^{p,q}$ the conjugacy of $\check{B}^{solv}$ to upper-triangular matrices can be achieved in $\mbox{CO}(p,q)$. Consequently, on ${\bf R}^{p+1,q+1}$, the conjugacy can be achieved in $P$, since $\check{B}^{solv}$ preserves $\ell_{0}$, and we can take the last basis vector to be a null vector not orthogonal to $\ell_{0}$.
@@
Fix a basis $X_{1},\ldots,X_{p^{\prime}}$ of ${\mathfrak{a}}_{H}$. Let $\check{X}_{1}$ be a lift of $X_{1}$ to $\check{{\mathfrak{b}}}^{solv}$. The semisimple component of $\check{X}_{1}$ in its Jordan decomposition in ${\mathfrak{g}}$ is nontrivial, so it is contained in $\check{{\mathfrak{b}}}^{solv}$ by algebraicity of the latter. Thus write $\check{X}_{1}=A_{1}+Z_{1}$ where $A_{1}$ is ${\bf R}$-semisimple in ${\mathfrak{g}}$ and $Z_{1}\in\ker r$. The subalgebra $r^{-1}({\mathfrak{a}}_{H})$ admits an ad $\mbox{ad }A_{1}$-invariant subspace decomposition $\check{{\mathfrak{a}}}_{1}\ltimes\ker r$. Now $\check{{\mathfrak{a}}}_{1}$ maps bijectively under $r$ to ${\mathfrak{a}}_{H}$, which is abelian, so it commutes with $A_{1}$.
Let $A_{2}$ be a lift of $X_{2}$ to $\check{{\mathfrak{a}}}_{1}$. The semisimple Jordan component of $A_{2}$ commutes with $A_{1}$, has the same projection to ${\mathfrak{a}}_{H}$ as $A_{2}$, and is contained in $\check{{\mathfrak{b}}}^{solv}$ by algebraicity. So we can assume $A_{2}$ is semisimple. The 00-eigenspace of ad $\mbox{ad }A_{1}$, denote it $E_{0}(A_{1})$, is ad $\mbox{ad }A_{2}$-invariant. There is thus an ad $\mbox{ad }A_{2}$-invariant complement to $\ker r$ in $r^{-1}({\mathfrak{a}}_{H})$, call it $\check{{\mathfrak{a}}}_{2}$, which is contained in $E_{0}(A_{1})$ and maps under $r$ onto ${\mathfrak{a}}_{H}$. This $\check{{\mathfrak{a}}}_{2}$ centralizes $A_{1}$ and $A_{2}$. Continuing for $i^{\prime}=3,\ldots,p^{\prime}$ results in an abelian subalgebra $\check{{\mathfrak{a}}}_{p^{\prime}}\subset\check{{\mathfrak{b}}}^{solv}$ mapping isomorphically under $r$ to ${\mathfrak{a}}_{H}$ and comprising ${\bf R}$-semisimple elements of ${\mathfrak{g}}$. Such a subalgebra is contained in a maximal abelian ${\bf R}$-split subalgebra of ${\mathfrak{p}}$, which is in turn conjugate in $P$ to the desired $\check{{\mathfrak{a}}}$ contained in our chosen diagonal subalgebra ${\mathfrak{a}}_{G}$.
@@ 4.2. Consequences of equivariance for stabilizer and isotropy
The facts that ${\mathfrak{p}}$ is a subalgebra of $\mathfrak{g}$ and that ad $\mbox{ad }{\mathfrak{p}}$ preserves a $(p,q)$-scalar product on $\mathfrak{g}/{\mathfrak{p}}$ up to scale have the following consequences via the embedding.
@@
    $\mathfrak{h}^{0}$ is ad $\mbox{ad }{\mathfrak{b}}_{H}$-invariant
@@
    ad $\mbox{ad }{\mathfrak{b}}_{H}$ preserves a symmetric bilinear form on $\mathfrak{h}/\mathfrak{h}^{0}$ up to scaling by $-2\rho(\beta_{1})$.
@@
If $\mathfrak{h}=\mathfrak{h}^{0}$, then (11) and injectivity of $\iota$ imply the image of $\iota$ is a semisimple subalgebra of ${\mathfrak{p}}$. Now $\iota(\mathfrak{h})=\iota_{\hat{x}}(\mathfrak{h})\subseteq{\mathfrak{p}}$ for $\hat{x}$ given by the Embedding Theorem, and it is an isotropy algebra at the point $x=\pi(\hat{x})$. The Thurston Stability Theorem [35, Thm 3] implies that $\mathfrak{h}$ is linearizable at $x$, which means $\iota(\mathfrak{h})$ is conjugate by an element $g\in P_{+}$ into ${\mathfrak{q}}$. Semisimplicity implies the image is in ${\mathfrak{q}}_{ss}$.
@@
These facts all follow easily from the ad $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$ in Proposition 4.3 (1), together with bracket relations in $\mathfrak{h}$:
@@
    For any nonzero $X\in\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$, the image $[\mathfrak{h}_{\alpha_{j^{\prime}}},X]=\mathfrak{h}_{-\alpha_{i^{\prime}}}$.
@@
    If $p^{\prime}>1$ and $\bar{\mathfrak{h}}_{-\alpha_{j^{\prime}}}\neq 0$ for $j^{\prime}\neq i^{\prime}$, then $[s]$ has an isotropic subspace of dimension greater than $12p^{\prime}-$ 1, unless $p^{\prime}=2$, $\rho(\beta_{1})=\alpha_{1}+\alpha_{2}$, and $[s]$ has an isotropic subspace of dimension $3$.
@@
    If $p^{\prime}>1$ and $\mathfrak{h}_{-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$, then $[s]$ has an isotropic subspace of dimension at least $12p^{\prime}-$ 1. If the maximal isotropic subspace for $[s]$ is dimension $12p^{\prime}-$ 1, then
@@
As in Observation 3 of [29], the idea is that if $\alpha+\alpha^{\prime}$ does not equal the conformal factor $-\rho(2\beta_{1})$, then $\overline{\mathfrak{h}}_{\alpha}\perp\overline{\mathfrak{h}}_{\alpha^{\prime}}$ for $[s]$.
@@
[…]
all occur as weights on $\mathfrak{h}/\mathfrak{h}^{0}$. Any sum of two of them has the form $\sum_{k^{\prime}=1}^{p^{\prime}}-n_{k^{\prime}}\alpha_{k^{\prime}}$ where $n_{k^{\prime}}\in{\bf N}$, $2\leq\sum n_{k^{\prime}}\leq 4$, and $n_{i^{\prime}}+n_{j^{\prime}}\geq 2$. Each such sum is obtained in one or two ways—that is, it is the sum of one or two different unordered pairs of elements of $\Sigma$. Excluding at most two elements of $\Sigma$ yields a subset $\Sigma^{\prime}$ so that the sum of any two elements of $\Sigma^{\prime}$ does not equal $-\rho(2\beta_{1})$. The root spaces $\mathfrak{h}_{-\alpha_{k^{\prime}}-\alpha_{\ell^{\prime}}}$ with $k^{\prime}\neq\ell^{\prime}$ have dimension 2, while $\mathfrak{h}_{-2\alpha_{i^{\prime}}},\mathfrak{h}_{-2\alpha_{j^{\prime}}}$ are 1-dimensional; the spaces $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}},\overline{\mathfrak{h}}_{-\alpha_{j^{\prime}}}$ have dimension at least $1$. Thus $\oplus_{\sigma\in\Sigma^{\prime}}\overline{\mathfrak{h}}_{\sigma}$ is totally isotropic for $[s]$ and has dimension at least $64p^{\prime}-$ 6. If $p^{\prime}\geq 3$, then the latter dimension exceeds $12p^{\prime}-$ 1.

### v2

@@ 3. Embedding Theorem
$\iota$ $:$ $\hat{M}\rightarrow\textrm{Mon}({\mathcal{S}},{\bf V})$
@@
Let $\mathbb{U}=\mbox{Mon}({\mathcal{S}},{\bf V})/P$ and $\bar{\iota}:M\rightarrow\mathbb{U}$ the induced map, which is also ${\textrm{Aut}}(M,\omega)$-equivariant. Note that if $g\in{\textrm{Aut}}(M,\omega)$ stabilizes $x$, then $\psi(g)$ stablizes $\bar{\iota}(x)$. Although $\mathbb{U}$ is not generally an algebraic variety, the Rosenlicht stratification [30] (see also [17, 2.2]) yields a decomposition
@@
with ${\textrm{Aut}}(M,\omega)$ acting by precomposition via $\psi$ on the first factor and trivially on the second factor, commuting with the $P$-action on the product. Considering the map $\iota\times K:\hat{M}\rightarrow\mathbb{U}$ given by $)\hat{x}\mapsto(\iota_{\hat{x}},K(\hat{x}$)), the same proof with the Borel Density Theorem as above yields, for almost-every $x\in M$, not only the embedding of $(\mathcal{S},\psi(B))$ into $({\bf V},\rho(P))$ but also that K(x^)K($\hat${x}) is fixed by $\check{B}_{\hat{x}}$; the conclusion that $\check{B}_{\hat{x}}$ contains the isotropy of $B$ at $x$ with respect to $\hat{x}$ moreover remains true.
@@
[…]
@@ 4. Application in conformal geometry
The remainder of the paper is devoted to the proof of Theorem 1.2. The first and longer part of the proof is algebraic and culminates in Proposition 4.12. We set $G=\mbox{O}(p+1,q+1)$ and $H$ a group locally isomorphic to $\mbox{SU}(p^{\prime},q^{\prime})$, with $p\leq q$ and $p^{\prime}<q^{\prime}$ as in the theorem.
@@
If $p<q$ then the root system of ${\mathfrak{o}}(p+1,q+1)$ is of type $B_{p+1}$, while if $p=q$, then it is type $D_{p+1}$. If $p^{\prime}<q^{\prime}$, then the root system of $\mathfrak{su}(p^{\prime},q^{\prime})$ is of type $BC_{p^{\prime}}$, while if $p^{\prime}=q^{\prime}$, then it is of type $C_{p^{\prime}}$. It is always assumed that $p,p^{\prime}>0$.
If $p<q$, let $\beta_{1},\ldots,\beta_{p+1}\in{\mathfrak{a}}^{*}$ be such that the positive roots of $\mathfrak{g}$ are
@@
The root system is $B_{p+1}$. The set $\{\beta_{k}\}$ are called the positive short roots, and their root spaces are each of dimension qpq-p. The remaining two subsets are the long roots, such that the $\{\beta_{i}-\beta_{j}\}$ are the initial roots of a string of positive roots, while the $\{\beta_{i}+\beta_{j}\}$ are the terminal roots of a string of positive roots. Either of these types of roots have $1$-dimensional root spaces. These strings are such that for any $X\in{\mathfrak{g}}_{\beta_{i}-\beta_{j}}$ and $Y\in{\mathfrak{g}}_{\beta_{j}}$ both nonzero, the brackets $X,Y]\in{\mathfrak{g}}_{\beta_{i}}$ and $X,Y],Y]\in{\mathfrak{g}}_{\beta_{i}+\beta_{j}}$ are nonzero. Also for nonzero $X\in{\mathfrak{g}}_{\beta_{i}-\beta_{j}}$ and $Y\in{\mathfrak{g}}_{\beta_{j}+\beta_{k}}$, the bracket $[X,Y]$ spans ${\mathfrak{g}}_{\beta_{i}+\beta_{k}}$. The weights of the $P$-representation on $\mathfrak{g}/{\mathfrak{p}}$ are
@@
in particular, the conformal factor is given by $2\beta_{1}$.
If $p=q$, we choose similiarly $\beta_{1},\ldots,\beta_{p+1}\in{\mathfrak{a}}^{*}$ such that the positive roots of $\mathfrak{g}$ are
$\Delta^{+}_{G}=\{\beta_{i}-\beta_{j}:i<j\}\cup\{\beta_{i}+\beta_{j}:i\neq j\}$
The root system is $D_{p+1}$. All these positive roots are of the same length and have $1$-dimensional root spaces. The weights of the representation $\mathfrak{g}/{\mathfrak{p}}$ are the same as in the case $p<q$, except that $\beta_{1}$ is not a weight; the conformal factor is still $2\beta_{1}$.
@@
For the first application of the Embedding Theorem, we consider a Borel subgroup $B_{H}<H$. Note that $B_{H}$ is solvable, algebraic, and discompact. As $B_{H}$ is amenable and $M$ is compact, there is a $B_{H}$-invariant probability measure $\mu$ on $M$. Let $\check{B}=\check{B}_{\hat{x}}<\mbox{Ad}_{\mathfrak{g}}P$ be the algebraic subgroup given by Theorem 3.1 applied to $B_{H}$ acting on the adjoint tractor bundle endowed with Čap’s connection as in Section 1.1. The parallel sections correspond to infinitesimal automorphisms, and we restrict to $\mathcal{S}=\mathfrak{h}$. Note that this version of the embedding theorem is exactly the one of [3].
@@ 4.1. Alignment of the intertwining epimorphism
The Lie algebra $\check{{\mathfrak{b}}}$ is a subalgebra of $\mathfrak{p}$}, with adjoint action preserving the image $\iota(\mathfrak{h})$, such that the restriction to this image is a surjection onto $\iota\circ\mbox{ad}_{\mathfrak{h}}{\mathfrak{b}}_{H}\circ\iota^{-1}\cong{\mathfrak{b}}_{H}$. We will denote by $r$ the resulting Lie algebra epimorphism $\check{{\mathfrak{b}}}\rightarrow{\mathfrak{b}}_{H}$. We can focus on the solvable radical $\check{{\mathfrak{b}}}^{solv}$, as any semisimple Levi factor is in the kernel of $r$. Because it is algebraic, $\check{{\mathfrak{b}}}^{solv}$ is invariant under Jordan decomposition (see [25, Thm 4.3.3]). Since $\mbox{Ad }B_{H}$ has no elliptic elements, any elliptic elements of $\check{{\mathfrak{b}}}^{solv}$ are in the kernel of $r$; more precisely, they belong to a compact torus of $\ker r$, which is a direct factor of $\check{{\mathfrak{b}}}^{solv}$ (see [25, Thm 4.4.7, Cor 4.4.12]), so we will simply exclude this factor and assume that all elements of $\check{{\mathfrak{b}}}^{solv}$ have trivial elliptic components in their Jordan decomposition.
@@
Lie’s Theorem ensures that in $V^{p,q}$ there is a simultaneous eigenvector $v_{1}$ or a $2$-dimensional irreducible subspace for $\check{B}^{solv}$. In the latter case, the restriction of $\check{B}^{solv}$ splits into a nontrivial elliptic and a scalar action. By the assumption that all elements of $\check{B}^{solv}$ have trivial elliptic Jordan components, this possibility is excluded. If the eigenvector $v_{1}$ is timelike or spacelike, then the eigenvalues on it are $\pm 1$. In this case, apply Lie’s Theorem again to $v_{1}^{\perp}$, to conclude that it contains an isotropic simultaneous eigenvector, or $\check{{\mathfrak{b}}}^{solv}$ is trivial on $V^{p,q}$. Proceeding inductively yields a null-orthonormal basis in which $\check{B}^{solv}$ is upper-triangular. It follows that on $V^{p,q}$ the conjugacy of $\check{B}^{solv}$ to upper-triangular matrices can be achieved in $\mbox{CO}(p,q)$. Consequently, on ${\bf R}^{p+1,q+1}$, the conjugacy can be achieved in $P$, since $\check{B}^{solv}$ preserves $\ell_{0}$, and we can take the last basis vector to be a null vector not orthogonal to $\ell_{0}$.
@@
Fix a basis $X_{1},\ldots,X_{p^{\prime}}$ of ${\mathfrak{a}}_{H}$. Let $\check{X}_{1}$ be a lift of $X_{1}$ to $\check{{\mathfrak{b}}}^{solv}$. The semisimple component of $\check{X}_{1}$ in its Jordan decomposition in $\mathfrak{g}$} is nontrivial, so it is contained in $\check{{\mathfrak{b}}}^{solv}$ by algebraicity of the latter. Thus write $\check{X}_{1}=A_{1}+Z_{1}$ where $A_{1}$ is $\bf R$}-semisimple in $\mathfrak{g}$} and $Z_{1}\in\ker r$. The subalgebra $r^{-1}({\mathfrak{a}}_{H})$ admits an $\mbox{ad }A_{1}$-invariant subspace decomposition $\check{{\mathfrak{a}}}_{1}\ltimes\ker r$. Now $\check{{\mathfrak{a}}}_{1}$ maps bijectively under $r$ to ${\mathfrak{a}}_{H}$, which is abelian, so it commutes with $A_{1}$.
Let $A_{2}$ be a lift of $X_{2}$ to $\check{{\mathfrak{a}}}_{1}$. The semisimple Jordan component of $A_{2}$ commutes with $A_{1}$, has the same projection to ${\mathfrak{a}}_{H}$ as $A_{2}$, and is contained in $\check{{\mathfrak{b}}}^{solv}$ by algebraicity. So we can assume $A_{2}$ is semisimple. The 0-eigenspace of $\mbox{ad }A_{1}$, denote it $E_{0}(A_{1})$, is $\mbox{ad }A_{2}$-invariant. There is thus an $\mbox{ad }A_{2}$-invariant complement to $\ker r$ in $r^{-1}({\mathfrak{a}}_{H})$, call it $\check{{\mathfrak{a}}}_{2}$, which is contained in $E_{0}(A_{1})$ and maps under $r$ onto ${\mathfrak{a}}_{H}$. This $\check{{\mathfrak{a}}}_{2}$ centralizes $A_{1}$ and $A_{2}$. Continuing for $i^{\prime}=3,\ldots,p^{\prime}$ results in an abelian subalgebra $\check{{\mathfrak{a}}}_{p^{\prime}}\subset\check{{\mathfrak{b}}}^{solv}$ mapping isomorphically under $r$ to ${\mathfrak{a}}_{H}$ and comprising $\bf R$}-semisimple elements of $\mathfrak{g}$}. Such a subalgebra is contained in a maximal abelian $\bf R$}-split subalgebra of $\mathfrak{p}$}, which is in turn conjugate in $P$ to the desired $\check{{\mathfrak{a}}}$ contained in our chosen diagonal subalgebra ${\mathfrak{a}}_{G}$.
@@ 4.2. Consequences of equivariance for stabilizer and isotropy
The facts that $\mathfrak{p}$} is a subalgebra of $\mathfrak{g}$ and that $\mbox{ad }{\mathfrak{p}}$ preserves a $(p,q)$-scalar product on $\mathfrak{g}/{\mathfrak{p}}$ up to scale have the following consequences via the embedding.
@@
    $\mathfrak{h}^{0}$ is $\mbox{ad }{\mathfrak{b}}_{H}$-invariant
@@
    $\mbox{ad }{\mathfrak{b}}_{H}$ preserves a symmetric bilinear form on $\mathfrak{h}/\mathfrak{h}^{0}$ up to scaling by $2\rho(\beta_{1})$.
@@
[…]
@@
These facts all follow easily from the $\mbox{ad }{\mathfrak{b}}_{H}$-invariance of $\mathfrak{h}^{0}$ in Proposition 4.3 (1), together with bracket relations in $\mathfrak{h}$:
@@
    For any nonzero $X\in\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$, the image $\mathfrak{h}_{\alpha_{j^{\prime}}},X]=\mathfrak{h}_{-\alpha_{i^{\prime}}}$.
@@
    If $p^{\prime}>1$ and $\bar{\mathfrak{h}}_{-\alpha_{j^{\prime}}}\neq 0$ for $j^{\prime}\neq i^{\prime}$, then $[s]$ has an isotropic subspace of dimension greater than $2p^{\prime}-1$, unless $p^{\prime}=2$, $\rho(\beta_{1})=\alpha_{1}+\alpha_{2}$, and $[s]$ has an isotropic subspace of dimension $3$.
@@
    If $p^{\prime}>1$ and $\mathfrak{h}_{-\alpha_{j^{\prime}}}\subset\mathfrak{h}^{0}$ for all $j^{\prime}\neq i^{\prime}$, then $[s]$ has an isotropic subspace of dimension at least $2p^{\prime}-1$. If the maximal isotropic subspace for $[s]$ is dimension $2p^{\prime}-1$, then
@@
As in Observation 3 of [29], the idea is that if $\alpha+\alpha^{\prime}$ does not equal the conformal factor $\rho(2\beta_{1})$, then $\overline{\mathfrak{h}}_{\alpha}\perp\overline{\mathfrak{h}}_{\alpha^{\prime}}$ for $[s]$.
@@
$\Sigma=\{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}},-\alpha_{j^{\prime}}-\alpha_{k^{\prime}},-\alpha_{i^{\prime}},-\alpha_{j^{\prime}}\ :\ 1\leq k^{\prime}\leq p^{\prime}\}$
all occur as weights on $\mathfrak{h}/\mathfrak{h}^{0}$. Any sum of two of them has the form $\sum_{k^{\prime}=1}^{p^{\prime}}-n_{k^{\prime}}\alpha_{k^{\prime}}$ where $n_{k^{\prime}}\in{\bf N}$, $2\leq\sum n_{k^{\prime}}\leq 4$, and $n_{i^{\prime}}+n_{j^{\prime}}\geq 2$. Each such sum is obtained in one or two ways—that is, it is the sum of one or two different unordered pairs of elements of $\Sigma$. Excluding at most two elements of $\Sigma$ yields a subset $\Sigma^{\prime}$ so that the sum of any two elements of $\Sigma^{\prime}$ does not equal $\rho(2\beta_{1})$. The root spaces $\mathfrak{h}_{-\alpha_{k^{\prime}}-\alpha_{\ell^{\prime}}}$ with $k^{\prime}\neq\ell^{\prime}$ have dimension 2, while $\mathfrak{h}_{-2\alpha_{i^{\prime}}},\mathfrak{h}_{-2\alpha_{j^{\prime}}}$ are 1-dimensional; the spaces $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}},\overline{\mathfrak{h}}_{-\alpha_{j^{\prime}}}$ have dimension at least $1$. Thus $\oplus_{\sigma\in\Sigma^{\prime}}\overline{\mathfrak{h}}_{\sigma}$ is totally isotropic for $[s]$ and has dimension at least $4p^{\prime}-6$. If $p^{\prime}\geq 3$, then the latter dimension exceeds $2p^{\prime}-1$.